# Torsion of $\mathbb{Q}$-curves over number fields of small odd prime degree

### v1

@@ 4.2. Possible odd orders of points
Proof.   If $E$ is such a $\mathbb{Q}$-curve, then by Corollary 2.6 there would be a rational elliptic curve with a $35$-isogeny over an extension of degree $5$ or $7$, which is impossible by [17, Theorem 1.3]. $\Box$
@@
Proof.   Suppose that there is a $\mathbb{Q}$-curve $E$ over a number field $K$ of degree $5$ or $7$ with a point of order $15$ over an extension of degree $5$ or $7$. Then we can copy the proof of Proposition 3.3 to conclude that we must do the following for each of the four curves $E^{\prime\prime}$ from the class 1600.i:
@@
  • In each of the subgroups $H$ from the list, find a matrix $A$ such that at least one matrix among $A\pmod{3}$, $A\pmod{5}$ doesn’t have $1$ as an eigenvalue, and at least one matrix among $A\pmod{3}$, $A\pmod{5}$ doesn’t have $-1$ as an eigenvalue.
@@
For each of the four curves in the isogeny class, the mod $21$ image has seven subgroups of index $7$. Each of those contains an element which does not have neither $1$ nor $-1$ as an eigenvalue modulo $7$, which completes the proof. $\Box$
4.3. Combining the two-torsion with the odd part
Note that for any $k>2$, the entire $k$-torsion $E[k]$ cannot be defined over an odd degree number field $K$, as $K(\zeta_{k})$ is contained in the field $K(E[k])$ and $\zeta_{k}$ is not contained in $K$.
A there are no points of order $16$ this means that the possible $2$-torsion over $K$ is either trivial or of the form $\mathbb{Z}/2^{k}\mathbb{Z}$ or $\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2^{k}\mathbb{Z}$ for $k\in\{1,2,3\}$.
@@ 4.3. Combining the two-torsion with the odd part
Proof.   If $E$ was such a $\mathbb{Q}$-curve, then there would be a rational elliptic curve $E^{\prime\prime}$ with a $14$-isogeny over an extension of degree $5$ or $7$, which is impossible by [17, Theorem 1.3], unless $E^{\prime\prime}$, and hence also $E$, has CM. However, by [3, 5.5 and 5.7], this is also not possible. $\Box$
@@
There are no $\mathbb{Q}$-curves over number fields of degree $5$ or $7$ with a point of order $18$ or $24$, or independent points of order $2$ and order $12$. Furthermore, there are no $\mathbb{Q}$-curves over number fields of degree $7$ with a point of order $20$, or independent points of order $2$ and $10$.
Proof.   For the first part, note that $5$ and $7$ do not divide $|\textnormal{GL}_{2}(\mathbb{Z}/3\mathbb{Z})|$ and $|\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})|$.
Now if $E$ is a $\mathbb{Q}$-curve over a number field of degree $5$ or $7$ with the mentioned torsion group, we can assume that there is a $2^{k}\cdot 3^{\ell}$-isogeny to an elliptic curve $E^{\prime}$ with rational $j$-invariant. Let $E^{\prime\prime}/\mathbb{Q}$ be the rational twist of $E^{\prime}$, and let $E^{\prime\prime\prime}$ be the corresponding twist of $E$. Then there is A $2^{k}3^{\ell}$-isogeny $E^{\prime\prime}\to E^{\prime\prime\prime}$ defined over $K$. Since $5$ and $7$ do not divide $|\textnormal{GL}_{2}(\mathbb{Z}/3\mathbb{Z})|$ and $|\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})|$, it follows that this isogeny is in fact defined over $\mathbb{Q}$, so $E^{\prime\prime\prime}$ is also rational. We conclude that $E$ has rational $j$-invariant. By the results in [6] for $K$ of degree $5$, and [16] for $K$ of degree $7$, and using [9, Theorem 5.1.2], we reach a contradiction.
Applying the exact same argument and noting that $7$ does not divide $|\textnormal{GL}_{2}(\mathbb{Z}/5\mathbb{Z})|$ and $|\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})|$ proves the second part. $\Box$
The proof of Theorem 1.10 is now complete.
@@
Proof.   Let $E$ be a $\mathbb{Q}$-curve over a number field $K$ of degree $5$ with two independent points of order $2$ and $10$ or with a point of order $20$. In both of these cases, $E$ is $K$-isogenous to an elliptic curve $E_{1}$ with a $20$-isogeny. Note that $E_{1}$ is also a $\mathbb{Q}$-curve.
Now let $E^{\prime}$ be an elliptic curve with rational $j$-invariant $K$-isogenous to $E_{1}$. Then, by Corollary 2.6, $E^{\prime}$ has a $5$-isogeny over $K$, and either has a $4$-isogeny or two independent $2$-isogenies over $K$. The same is true for a rational twist $E^{\prime\prime}$ of $E^{\prime}$.
But then we have a rational elliptic curve with a $5$-isogeny over $K$ and either a $4$-isogeny or two independent $2$-isogenies. Since $5\nmid|\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})|$, these $2$-isogenies or the $4$-isogeny is defined over $\mathbb{Q}$. By Proposition 2.8, there is also a $5$-isogeny over $\mathbb{Q}$.
However, this means that $E^{\prime\prime}$ either has a rational $20$-isogeny or a rational $2$-isogeny and a $10$-isogeny, in which case it is $2$-isogenous over $\mathbb{Q}$ to an elliptic curve with a rational $20$-isogeny. which contradicts [17, Theorem 1.3.]. $\Box$
@@
Proof.   Let $E$ be a $\mathbb{Q}$-curve over a quintic number field $K$ with a point $P$ of order $50$. Then there is an isogenous elliptic curve $E^{\prime}$ with rational $j$-invariant whose Galois image is conjugate to a subgroup of all matrices of the form $\begin{bmatrix}1&\frac{50}{\alpha}\cdot k\\ \alpha\cdot j&z\end{bmatrix}$ for some $\alpha$ dividing $50$. Without loss of generality, we may assume that the isogeny is in fact a $5^{k}$-isogeny for some $k$, and then $\alpha$ is either $2$, $10$ or $50$.
If $\alpha=50$, we reach a contradiction as an elliptic curve with rational $j$-invariant cannot have a rational point of order $50$.
If $\alpha\in\{2,10\}$, $\rho_{E^{\prime},25}(G_{K})$ has cardinality dividing $5^{3}\cdot 4$. If we denote by $E^{\prime\prime}$ a rational twist of $E^{\prime}$, then $\rho_{E^{\prime\prime},25}(G_{K})$ has cardinality dividing $5^{4}\cdot 8$. Also, $E^{\prime\prime}$ has a rational $5$-isogeny by Proposition 2.8. By [8, Theorem 2], if none of the curves in the $\mathbb{Q}$-isogeny class of $E^{\prime\prime}$ has two independent $5$-isogenies, then $\rho_{E^{\prime\prime},5^{\infty}}(G_{\mathbb{Q}})$ contains a Sylow pro-$p$ subgroup. However, the Sylow pro-$p$ subgroup when viewed modulo $25$ has $5^{5}$ elements, which does not divide $5^{4}\cdot 4$- Thus, one of the curves in the isogeny class of $E^{\prime\prime}$ has two independent rational $5$-isogenies. Since $E^{\prime\prime}$ also has a rational $2$-isogeny coming from $E$, it follows that there is a rational elliptic curve with a $50$-isogeny, which is a contradiction. $\Box$
This completes the proof of Theorem 1.9.
@@ 5. Quadratic torsion of elliptic curves isogenous to those with rational $j$-invariant
In this Section, we prove Theorem 1.13
@@
Proof.   Suppose that $E(K)$ has torsion $\mathbb{Z}/14\mathbb{Z}$. Let $E^{\prime}$ be the elliptic curve with rational $j$-invariant isogenous to $E$. Applying a quadratic twist, one may assume that $E^{\prime}$ itself is defined over $K$. Then, by [4, Lemma A.4], there is a quadratic twist $E^{\prime\prime}$ of $E^{\prime}$ such that there is an isogeny $E\to E^{\prime\prime}$ defined over $K$ itself. By Proposition 2.5, it follows that $E^{\prime\prime}$ has a $14$-isogeny over $K$, which is not a possible isogeny degree of elliptic curves with rational $j$-invariant over quadratic fields by [20].
Suppose that $E(K)$ has torsion $\mathbb{Z}/18\mathbb{Z}$. By the results in [1, Section 4.6.], it follows that $E$ is $2$-isogenous to its Galois conjugate. If $E$ is isogenous to an elliptic curve with rational $j$-invariant then, using notation from [4, Appendix], the $j$-invariant of $E$ belongs to a rational class $\mathcal{Q}$. However, the degree of an isogeny from $E$ to its conjugate then has to be a square by [4, Corollary A.12] and it cannot be $2$Box
Proof of Theorem 1.13. For $p>2$ this follows from Corollary 1.12, as elliptic curves over odd prime degree number fields $K$ are exactly those isogenous to elliptic curves with rational $j$-invariant.
For $p=2$, we know from the list of all possible torsion subgroups over quadratic fields by the results from Kamienny [10], and Kenku and Momose [11]. Najman [16, Theorem 2] found the list of all torsion subgroups of rational elliptic curves over quadratic extensions, and Gužvić [9, Theorem 5.1.2.] found the list of all torsion subgroups of elliptic curves with rational $j$-invariant over quadratic extensions.
It follows from these results that $\mathbb{Z}/14\mathbb{Z}$ and $\mathbb{Z}/18\mathbb{Z}$ are the only torsion subgroups that appear for some elliptic curves over quadratic fields and do not appear for elliptic curves with rational $j$-invariant. By Proposition 5.1, they also do not appear for elliptic curves isogenous to those with rational $j$-invariant, which completes the proof. ∎

### v2

@@ 4.2. Possible odd orders of points
Proof.   If $E$ is such a $\mathbb{Q}$-curve, then there exists an elliptic curve with rational $j$-invariant in the isogeny class with a $35$-isogeny over an extension of degree $5$ or $7$, which is impossible by [16, Theorem 1.3]. $\Box$
@@
Proof.   Suppose that there is a $\mathbb{Q}$-curve $E$ over a number field $K$ of degree $5$ or $7$ with a point of order $15$ over an extension of degree $5$ or $7$. Then an elliptic curve with rational $j$-invariant in the isogeny class has a quadratic twist which has a rational $15$-isogeny or a $21$-isogeny. We can now copy the proof of Proposition˜3.3 to conclude that we must do the following for each of the four curves $E^{\prime\prime}$ from the class 1600.i:
@@
  • In each of the subgroups $H$ from the list, find a matrix $A$ such that at least one matrix among $A\pmod{3}$, $A\pmod{5}$ doesn’t have $1$ as an eigenvalue, and at least one matrix among $A\pmod{3}$, $A\pmod{5}$ doesn’t have 1-1 as an eigenvalue.
@@
For each of the four curves in the isogeny class, the mod $21$ image has seven subgroups of index $7$. Each of those contains an element which does not have neither $1$ nor 1-1 as an eigenvalue modulo $7$, which completes the proof. $\Box$
4.3. Combining the $2^{k}$-torsion with the odd part
Note that for any $n>2$, the entire $n$-torsion $E[n]$ cannot be defined over an odd degree number field $K$, as $K(\zeta_{n})$ is contained in the field $K(E[n])$ and $\zeta_{n}$ is not contained in $K$.
As there are no points of order $16$ this means that the possible $2^{k}$-torsion over $K$ is either trivial or of the form $\mathbb{Z}/2^{k}\mathbb{Z}$ or $\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2^{k}\mathbb{Z}$ for $k\in\{1,2,3\}$.
@@ 4.3. Combining the $2^{k}$-torsion with the odd part
Proof.   If $E$ was such a $\mathbb{Q}$-curve, then there would be a rational elliptic curve $E^{\prime\prime}$ with a $14$-isogeny over an extension of degree $5$ or $7$, which is impossible by [16, Theorem 1.3], unless $E^{\prime\prime}$, and hence also $E$, has CM. However, by [3, 5.5 and 5.7], this is also not possible. $\Box$
@@
There are no $\mathbb{Q}$-curves over number fields of degree $5$ or $7$ with a point of order $18$ or $24$, or independent points of order $2$ and order $12$.
Proof.   If $E$ is a $\mathbb{Q}$-curve over a number field of degree $5$ or $7$ with the mentioned torsion group, we can assume that there is a $2^{k}\cdot 3^{\ell}$-isogeny to an elliptic curve $E^{\prime}$ with rational $j$-invariant. Let $E^{\prime\prime}/\mathbb{Q}$ be the rational twist of $E^{\prime}$, and let $E^{\prime\prime\prime}$ be the corresponding twist of $E$. Then there is a $2^{k}3^{\ell}$-isogeny $E^{\prime\prime}\to E^{\prime\prime\prime}$ defined over $K$. Since $5$ and $7$ do not divide $\textnormal{GL}_{2}(\mathbb{Z}/3\mathbb{Z})|$ and $\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})$, it follows that this isogeny is in fact defined over $\mathbb{Q}$, so $E^{\prime\prime\prime}$ is also rational. We conclude that $E$ has rational $j$-invariant. By the results in [7] for $K$ of degree $5$, and [15] for $K$ of degree $7$, and using [9, Theorem 5.1.2], we reach a contradiction. $\Box$
Lemma 4.10.
There are no $\mathbb{Q}$-curves over number fields of degree $7$ with a point of order $20$ or independent points of order $2$ and $10.$
Proof.   We apply the same argument as in the proof of the previous lemma but with an isogeny of degree $2^{k}5^{\ell}$, noting that $7$ does not divide $\textnormal{GL}_{2}(\mathbb{Z}/5\mathbb{Z})$ and $\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})$. We obtain an elliptic curve with rational $j$-invariant and it cannot have a point of order $20$ or independent points of order $2$ and $10$ by [9, Theorem 5.1.2]. $\Box$
The proof of Theorem˜1.11 is now complete, i.e. the possible torsion subgroups of elliptic curves over septic fields have all been determined..
@@
Proof.   Let $E$ be a $\mathbb{Q}$-curve over a number field $K$ of degree $5$ with two independent points of order $2$ and $10$ or with a point of order $20$. In both of these cases, $E$ is $K$-isogenous to an elliptic curve $E_{1}$ with a $20$-isogeny. The elliptic curve $E_{1}$ is $2^{a}\cdot 5^{b}$-isogenous to an elliptic curve $E^{\prime}$ with rational $j$-invariant. We may factor this $2^{a}\cdot 5^{b}$-isogeny as a $5^{b}$-isogeny $E\to E_{2}$ and a $2^{a}$-isogeny $E_{2}\to E^{\prime}$. Since $5$ does not divide $\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})$, it follows that $E_{2}$ also has rational $j$-invariant, so we may assume that $a=0$ and that $E_{1}$ is $5^{b}$-isogenous to $E^{\prime}$.
However, $E^{\prime}$ has a $4$-isogeny over $K$ since $E_{1}$ has a $4$-isogeny. Since it also has a $5$-isogeny, it has a $20$-isogeny. But $20$ is not a possible isogeny degree of an elliptic curve with rational $j$-invariant, so we have reached a contradiction. $\Box$
@@
Proof.   Let $E$ be a $\mathbb{Q}$-curve over a quintic number field $K$ with a point $P$ of order $50$. Then there is an isogenous elliptic curve $E^{\prime}$ with rational $j$-invariant. As in the previous proof, we may assume that the isogeny $E\to E^{\prime}$ is a $5^{b}$-isogeny for some $b\geq 1$. If $b\geq 3$, then $E^{\prime}$ is an elliptic curve with rational $j$-invariant and a $250$-isogeny over a quintic extension, which is impossible by [16, Theorem 1.3]. If the isogeny $E\to E^{\prime}$ does not factor through the $50$-isogeny whose kernel is the point of order $50$, then $E^{\prime}$ again has a $2\cdot 5^{2+b}$-isogeny which is impossible. Thus, $b\leq 2$ and the isogeny $E\to E^{\prime}$ factors through the isogeny whose kernel is the point of order $50$.
Suppose that $b=2$. We have the following diagram.
${\star}E{{E^{\prime}}}{\bullet}$
By Proposition˜2.8, we know that some quadratic twist $E^{\prime\prime}$ of $E^{\prime}$ has a rational $5$-isogeny, so either the elliptic curve marked with $\star$ or the elliptic curve marked with $\bullet$ has rational $j$-invariant. If the elliptic curve marked with $\bullet$ has rational $j$-invariant, then it has a $250$-isogeny over $K$, a contradiction. Thus, the elliptic curve marked with $\star$ has rational $j$-invariant. Thus we may actually assume that $b=1$ and treat the elliptic curve marked with $\star$ as $E^{\prime}$.
By Proposition˜2.5, $\rho_{E^{\prime},50}(G_{K})$ is contained in the subgroup of the group of all matrices of the form $\begin{bmatrix}1&5\cdot k\\ 10\cdot j&z\end{bmatrix}$. It follows that $\rho_{E^{\prime},25}(G_{K})$ divides $5^{3}\cdot 4$. Let $E^{\prime\prime}$ be a quadratic twist of $E^{\prime}$ which is defined over $\mathbb{Q}$. Then $\rho_{E^{\prime\prime},25}(G_{K})$ has cardinality dividing $5^{4}\cdot 8$. Also, $E^{\prime\prime}$ has a rational $5$-isogeny by Proposition˜2.8. By [8, Theorem 2], if none of the curves in the $\mathbb{Q}$-isogeny class of $E^{\prime\prime}$ has two independent $5$-isogenies, then $\rho_{E^{\prime\prime},5^{\infty}}(G_{\mathbb{Q}})$ contains a Sylow pro-$p$ subgroup. However, the Sylow pro-$p$ subgroup when viewed modulo $25$ has $5^{5}$ elements, which does not divide $5^{4}\cdot 8$. Thus, one of the curves in the $\mathbb{Q}$-isogeny class of $E^{\prime\prime}$ has two independent rational $5$-isogenies. Since $E^{\prime\prime}$ also has a rational $2$-isogeny coming from $E$, it follows that the isogeny class a rational elliptic curve with a $50$-isogeny, which is a contradiction. $\Box$
This completes the proof of Theorem˜1.10, i.e. the possible torsion subgroups over quintic fields have all been determined.
@@ 5. Quadratic torsion of elliptic curves isogenous to those with rational $j$-invariant
In this section, we prove Theorem˜1.14.
@@
Proof.   Suppose that $E(K)$ has torsion $\mathbb{Z}/14\mathbb{Z}$. Let $E^{\prime}$ be the elliptic curve with rational $j$-invariant isogenous to $E$. Applying a quadratic twist, one may assume that $E^{\prime}$ itself is defined over $K$. Then, by [4, Lemma A.4], there is a quadratic twist $E^{\prime\prime}$ of $E^{\prime}$ such that there is an isogeny $E\to E^{\prime\prime}$ defined over $K$ itself. By Proposition˜2.5, it follows that $E^{\prime\prime}$ has a $14$-isogeny over $K$, which is not a possible isogeny degree of elliptic curves with rational $j$-invariant over quadratic fields by [19].
Suppose that $E(K)$ has torsion $\mathbb{Z}/18\mathbb{Z}$. By the results in [1, Section 4.6.], it follows that $E$ is $2$-isogenous to its Galois conjugate. If $E$ is isogenous to an elliptic curve with rational $j$-invariant then, using notation from [4, Appendix], the $j$-invariant of $E$ belongs to a rational class $\mathcal{Q}$. However, the degree of an isogeny from $E$ to its conjugate then has to be a square by [4, Corollary A.12] and it cannot be $2$. $\Box$
Proof of Theorem˜1.14. For $p>2$ this follows from Corollary˜1.13, as elliptic curves over odd prime degree number fields $K$ are exactly those isogenous to elliptic curves with rational $j$-invariant.
For $p=2$, we know from the list of all possible torsion subgroups over quadratic fields by the results from Kamienny [10], and Kenku and Momose [11]. Najman [15, Theorem 2] found the list of all torsion subgroups of rational elliptic curves over quadratic extensions, and Gužvić [9, Theorem 5.1.2.] found the list of all torsion subgroups of elliptic curves with rational $j$-invariant over quadratic extensions.
It follows from these results that $\mathbb{Z}/14\mathbb{Z}$ and $\mathbb{Z}/18\mathbb{Z}$ are the only torsion subgroups that appear for some elliptic curves over quadratic fields and do not appear for elliptic curves with rational $j$-invariant. By Proposition˜5.1, they also do not appear for elliptic curves isogenous to those with rational $j$-invariant, which completes the proof. ∎